The position of an object is given by x = A t + B t 2, where x is in metres and t is in seconds. 

  1. What are the units of A and B? [ANS A = m/s, B = m/s2] [Homogeneity analysis] 
  2. What is the acceleration as a function of time? [ANS = 2B m/s2] [DIFFERENTIATION twice] 
  3. What are the velocity and acceleration at t = 5.0 s? [ANS v = (A + 10B) m/s, a = 2B m/s2] 
  4. What is the velocity as a function of time if x = A t + B t -3 ? [ANS = A – 3B t -4] 

 


1
Expert's answer
2021-11-07T17:10:54-0500

1) The total units of x must be meters, so, any element (At and Bt2) must output meter. Therefore, since we have time and time squared, the units of A and B are m/s and m/s2.

2) "x'(t)=A+2Bt^{2-1}=v(t),a=x''(t)=v'(t)=2B."

3) Velocity is

"v(5)=A+2Bt=A+10B,\\
a(5)=2B."


4) Velocity is


"v(t)=A+B\u00b7(-3)\u00b7t^{-3-1}=A-3Bt^{-4}."